# Investigating Efficacy of Perplexity in Detecting LLM-Generated Code

### v1

@@ V-E $RQ_{\text{3}}$: Inference Efficiency
TABLE IX: The detection speed of Perplexity in different LLMs (seconds per sample)
@@
TABLE X: The detection speed of feature-based and pre-training-based methods in different LLMs (seconds per sample)
@@
Table IX and Table X present the inference speed of Perplexity and the other two classes of methods for code generated by four LLMs, respectively. For presentation clarity, we highlight notable results with blue backgrounds (sec/sample:$0-0.01$;$0.01-0.1$;$0.1-1$)
@@
Perplexity is not as fast as the feature-based and pre-training-based methods in all four LLMs. The AUC of Perplexity and the pre-training-based method for each LLM are as follows: GPT3.5 (0.076–0.014), GPT-4o (0.080–0.010), Gemini (0.078–0.010), and Llama (0.079–0.010). Perplexity calculates the perplexity of each token (a code snippet consists of a set of tokens) using another LLM, therefore, it is very time-consuming. When compared with the feature-based method, Perplexity is slower than Idialu: GPT3.5 (0.076–0.038), GPT-4o (0.080–0.042), Gemini (0.078–0.040), and Llama (0.079–0.040). However, Perplexity is faster than Bukhari: GPT3.5 (0.076–4.352), GPT-4o (0.080–4.312), Gemini (0.078–4.504), and Llama (0.079–4.794). Since Bukhari extracts n-gram features from abstract syntactic trees, it has a slow inference speed.
@@
When comparing the inference speed of methods in Perplexity, we observe that methods without perturbation are much faster than methods with perturbation. The inference speed of the worst-performing method without perturbation is less than 2 seconds per sample: GPT3.5 (1.265), GPT-4o (1.379), Gemini (1.288), and Llama (0.801). However, the inference speed of the best-performing method with perturbation is more than 4 seconds per sample: GPT3.5 (5.306), GPT-4o (5.619), Gemini (5.370), and Llama (4.879). Perplexity with perturbations introduces variations into the original sample to generate a list of perturbed samples and calculate the attributes such as log-rank of two types of samples. When comparing Perplexity methods with perturbation, NPR and DetectCodeGPT have a faster inference speed than DetectGPT. The average speed of NPR and DetectCodeGPT is around 10 seconds per sample: GPT3.5 (11.97–11.97), GPT-4o (12.92–12.92), Gemini (12.03–12.03), and Llama (9.685–9.685). NPR and DetectCodeGPT perform very closely because DetectCodeGPT is based on NPR. However, the average inference speed of DetectGPT is more than 100 seconds per sample: GPT3.5 (152.9), GPT-4o (147.6), Gemini (156.8), and Llama (169.6). The perturbation method of DetectGPT is complex, i.e., masking and fulfilling tokens by another LLM. The perturbation method of DetectCodeGPT is simple, i.e., adding spaces and new lines to the original tokens.
@@
$\textbf{RQ}_{\textbf{3}}$: Perplexity has a lower inference speed than the feature-based and pre-training-based methods. The inference speed of Perplexity without perturbation is faster than Perplexity with perturbation.
VI Implications
This section discusses implications based on the experimental results and the detection mechanism of Perplexity.
VI-A For Researchers
Our implications for researchers include both pros and cons of Perplexity.
@@ VI-A For Researchers
  • Perplexity is unsuitable for high-level programming languages. Perplexity performs well in low-level languages such as C/C++ (91%) but poorly in high-level languages such as Python (64%) and Ruby (73%), which implies that Perplexity has limited capability for high-level languages. High-level languages encourage a concise and clear code style. High-level languages such as Python provide many built-in operations for basic elements, e.g., list and dictionary; low-level languages have to manually manage array, pointer, etc. The programming patterns of high-level languages can be easier learned by a code generator than that of low-level languages. Moreover, high-level languages have more available third-party libraries than low-level languages. Developers are used to leveraging third-party libraries in their daily development. Since humans and LLMs use third-party libraries in generating code, detecting LLMgCode from HaCode using Perplexity in high-level languages is challenging.
@@
  • Perplexity is unsuitable for just-in-time detection. NPR and DetectCodeGPT are the best-performing methods in Perplexity. The inference speeds of NPR and DetectCodeGPT are all 10 seconds per sample. On the contrary, the inference speed of the pre-training-based method is less than 0.02 seconds per sample. The difference implies that Perplexity is not suitable for just-in-time detection. This can be explained that Perplexity applies an LLM to calculate the perplexity of each token. The calculation efficiency of LLMs is much lower than small pre-trained models such as CodeBERT. Moreover, NPR and DetectCodeGPT are perturbation-based methods. Such a method calculates the perplexity for one original sample and a collection of perturbed samples. As the number of perturbed samples increases, the perturbation-based methods get time-consuming.
@@
Although our large-scale investigation indicates that Perplexity has several shortcomings, it exhibits advantages in the following aspects.
  • Perplexity exhibits good interpretability. Perplexity tries to simulate code generation mechanisms of LLMs to calculate the perplexity. Therefore, the perplexity value of a code snippet is interpretable. A code snippet with a high perplexity indicates that it is disjointed, which might not be LLMgCode. The threshold of perplexity for detection can be adjusted. Although several supervised methods perform better than Perplexity, their mechanisms are complex and hard to interpret.
@@
  • Perplexity exhibits overall good generalization capability regarding LLMs. Perplexity achieves similar performance in detecting LLMgCode across four LLMs, with AUC variations of around 2%. This implies that Perplexity has a good generalization capability when detecting LLMgCode that is generated by different LLMs. The detection mechanism of Perplexity is similar to LLMs for code generation. Therefore, Perplexity is useful as long as LLMs are based on the auto-regressive generation mechanism (i.e., generating a token based on tokens generated before). The performance of supervised methods drops when the code snippets for training and testing are generated by different LLMs. Considering new powerful LLMs are emerging, Perplexity is the most promising method to address this issue.
  • Perplexity has the potential for fine-grained detection. Perplexity calculates the perplexity of each token in the code snippet. This implies that Perplexity can be applied for fine-grained detection. For instance, if a code snippet includes both LLMgCode and HaCode, Perplexity can calculate the perplexity for each line and detect these lines with low perplexity as LLMgCode. However, feature-based and pre-training-based methods only output a binary result, i.e., the code is either LLMgCode or HaCode.
In short, Perplexity is not effective enough but is the most promising method overall. Unlike those supervised methods that aim to learn characteristics of one specific type of code (e.g., GPT 3.5-generated code), Perplexity seeks to understand intrinsic differences between LLMgCode and HaCode. Therefore, Perplexity has a good generalization capacity. By continuously improving Perplexity on the target dataset, it is expected that the improved performance can be generalized. Moreover, Perplexity can be applied naturally to fine-grained detection (e.g., line level) when LLMgCode and HaCode are mixed, whereas other methods only provide a binary classification result. This enhances the applicability of Perplexity in complex real-world scenarios.
VI-B For Tool Development and Usages

### v2

@@ 5.4. $\text{RQ}_{\text{2}}$: Detection Speed
Table 8. The detection speed of Perplexity in terms of LLMs (seconds per sample)
@@
Table 9. The detection speed of feature- and pre-training-based methods in terms of LLMs (seconds per sample)
@@
Table 8 and Table 9 present the speed of the three classes of methods in detecting code generated by four LLMs. For presentation clarity, we highlight notable results in blue (sec/sample: $0-0.01$;$0.01-0.1$;$0.1-1$)
@@
Perplexity is not as fast as feature- and pre-training-based methods on all four LLMs. The AUC of Perplexity and the pre-training-based method for each LLM are as follows: GPT-3.5 (0.076–0.014), GPT-4o (0.080–0.010), Gemini-1.0 (0.078–0.010), and Llama-3.1 (0.079–0.010). Perplexity calculates the perplexity of each token (a code snippet consists of a set of tokens) using another LLM; therefore, it is very time-consuming. When compared with the feature-based method, Perplexity is slower than Idialu: GPT-3.5 (0.076–0.038), GPT-4o (0.080–0.042), Gemini-1.0 (0.078–0.040), and Llama-3.1 (0.079–0.040). However, Perplexity is faster than Bukhari: GPT-3.5 (0.076–4.352), GPT-4o (0.080–4.312), Gemini-1.0 (0.078–4.504), and Llama-3.1 (0.079–4.794). Since Bukhari extracts n-gram features from abstract syntactic trees, it has a slow detection speed.
@@
When comparing the detection speed of methods in Perplexity, we observe that methods without perturbation are much faster than methods with perturbation. The speed of the worst-performing method without perturbation is less than 2 seconds per sample: GPT-3.5 (1.265), GPT-4o (1.379), Gemini-1.0 (1.288), and Llama-3.1 (0.801). On the contrary, the speed of the best-performing method with perturbation is more than 4 seconds per sample: GPT-3.5 (5.306), GPT-4o (5.619), Gemini-1.0 (5.370), and Llama-3.1 (4.879). Perplexity with perturbations introduces variations into the original sample to generate a list of perturbed samples and calculate attributes such as log-rank of two types of samples, whereas Perplexity without perturbations only calculates attributes of the original sample. Therefore, Perplexity without perturbations performs much faster than Perplexity with perturbations. When comparing Perplexity methods with perturbation, NPR and DetectCodeGPT have a faster detection speed than DetectGPT. The average speed of NPR and DetectCodeGPT is around 10 seconds per sample: GPT-3.5 (11.97–11.97), GPT-4o (12.92–12.92), Gemini-1.0 (12.03–12.03), and Llama-3.1 (9.685–9.685). NPR and DetectCodeGPT achieve very close performance because DetectCodeGPT is based on NPR. However, the average speed of DetectGPT is more than 100 seconds per sample: GPT-3.5 (152.9), GPT-4o (147.6), Gemini-1.0 (156.8), and Llama-3.1 (169.6). The perturbation method of DetectGPT is complex, i.e., masking and completing tokens by another LLM. The perturbation method of DetectCodeGPT is simple, i.e., adding spaces and new lines to the original tokens.
@@
$\textbf{RQ}_{\textbf{2}}$: Perplexity exhibits a lower detection speed than the feature- and pre-training-based methods. The detection speed of Perplexity without perturbation is faster than Perplexity with perturbation.
5.5. $\text{RQ}_{\text{3}}$: Generalization Capability
Table 10 and Table 11 present the AUC of Perplexity and the other two classes of methods in detecting LLMgCode generated by four LLMs. The LLMs listed in the two tables are used to generate LLMgCode. The code LLMs listed in Table 10 are used to calculate perplexity values. For presentation clarity, we highlight notable results in blue (AUC:$90\%-100\%$;$80\%-90\%$;$70\%-80\%$).
Table 10. The AUC of Perplexity in terms of LLMs
LLM Code LLM without perturbation with perturbation
log-p Ent Rank L-Rank LRR DGPT NPR DCGPT
GPT-3.5 Incoder 51.89 46.86 39.93 63.01 72.28 49.68 73.10 72.78
Phi-1 75.44 41.93 66.19 75.87 45.27 52.79 81.10 81.64
StarCoder 56.14 54.82 55.86 61.09 57.55 47.53 70.86 70.03
CodeGen2 63.27 49.26 58.56 66.84 56.52 48.27 73.31 72.40
WizCoder 66.34 42.77 60.98 68.54 50.25 72.24 74.26 74.37
CodeLlama 59.41 53.78 60.53 65.79 60.39 41.48 71.52 71.28
Average 62.08 48.24 57.01 66.86 57.04 52.00 74.03 73.75
GPT-4o Incoder 55.29 44.88 44.15 63.16 63.12 54.10 74.89 74.71
Phi-1 72.89 38.02 63.79 72.56 44.10 58.14 80.52 80.88
StarCoder 56.04 52.69 53.07 58.90 53.02 51.63 71.25 70.44
CodeGen2 62.52 48.15 54.12 63.91 51.03 53.03 72.10 71.55
WizCoder 65.54 40.62 57.90 65.91 45.73 66.75 72.95 73.00
CodeLlama 60.70 50.93 56.88 64.68 54.86 48.01 71.13 70.93
Average 62.16 45.88 54.98 64.85 51.98 55.28 73.81 73.59
Gemini-1.0 Incoder 54.53 45.43 42.55 63.70 76.00 51.37 73.74 73.17
Phi-1 74.55 42.25 71.10 74.30 51.08 54.88 81.78 81.66
StarCoder 60.58 53.99 64.18 64.75 66.84 49.24 73.36 72.26
CodeGen2 67.39 48.34 65.84 69.68 65.25 50.68 76.23 74.98
WizCoder 69.21 44.70 64.43 67.99 50.52 58.07 74.36 74.15
CodeLlama 63.76 53.00 68.01 66.36 68.45 43.83 73.49 73.21
Average 65.00 47.95 62.68 67.80 63.02 51.35 75.49 74.90
Llama-3.1 Incoder 57.00 42.88 43.15 65.10 74.82 50.84 73.20 73.14
Phi-1 74.45 41.33 68.64 73.47 46.39 54.96 79.42 79.55
StarCoder 66.20 47.63 67.67 70.29 72.05 51.47 75.56 75.23
CodeGen2 71.42 42.79 66.51 73.05 66.92 49.79 77.42 76.95
WizCoder 74.69 38.59 68.99 76.39 56.50 67.27 79.08 78.74
CodeLlama 69.59 47.23 71.41 73.78 72.37 52.15 77.52 77.69
Average 68.89 43.41 64.39 72.01 64.84 54.41 77.03 76.88
Table 11. The AUC of feature- and pre-training-based methods in terms of LLMs
LLM Feature Pre-training
Idialu (Python) Bukhari (C) GPTSniffer
GPT-3.5 91.43 98.63 92.21
GPT-4o 83.16 98.49 90.30
Gemini-1.0 62.77 92.13 75.21
Llama-3.1 71.76 97.15 83.26
Perplexity has a better generalization capacity than the feature- and pre-training-based methods. The AUC of Perplexity for four LLMs is: GPT-3.5 (81.64%), GPT-4o (80.88%), Gemini-1.0 (81.78%), and Llama-3.1 (79.55%), with a range of 2%. The AUC of pre-training-based methods for each LLM is: GPT-3.5 (92.21%), GPT-4o (90.30%), Gemini-1.0 (75.21%), and Llama-3.1 (83.26%). The max-min difference is over 17%. These differences can be explained that pre-training-based methods are trained on the code generated by GPT-3.5, while Perplexity is LLM-independent. When it comes to feature-based methods, Perplexity outperforms Idialu and Bukhari. The AUC of Idialu for each LLM is as follows: GPT-3.5 (91.43%), GPT-4o (83.16%), Gemini-1.0 (62.77%), and Llama-3.1 (71.76%). The max-min difference is 29%. The AUC of Bukhari for each LLM is as follows: GPT-3.5 (98.63%), GPT-4o (98.49%), Gemini-1.0 (92.13%), and Llama-3.1 (97.15%), with a range of 7%. The max-min range of Bukhari is relatively small. Bukhari only works for C in which HaCode is more different from LLMgCode than other languages, making the detection method easier to detect LLMgCode from HaCode.
When comparing the generalization capability of methods in Perplexity, we observe that methods with perturbation have better generalization than methods without perturbation. Methods with perturbation achieve similar performance on the four LLMs, with the range of average AUC around 3%. For instance, the average AUC of DetectCodeGPT for the four LLMs is: GPT-3.5 (73.75%), GPT-4o (73.59%), Gemini-1.0 (74.90%), and Llama-3.1 (76.88%). In contrast, methods without perturbation do not exhibit such performance, with some methods fluctuating up to 10% on different LLMs. For instance, the average AUC of Rank for the four LLMs is as follows: GPT-3.5 (57.01%), GPT-4o (54.98%), Gemini-1.0 (62.68%), and Llama-3.1 (64.39%). Perplexity without perturbations are based on the absolute perplexity value of the original sample. The absolute perplexity is calculated based on the understanding of the code by the detection model. Therefore, the absolute perplexity varies across code generated by different LLMs, leading to fluctuations in the performance of Perplexity without perturbations. In contrast, Perplexity with perturbations are based on the perplexity discrepancy between the original and perturbed samples, making them more generalized.
$\textbf{RQ}_{\textbf{3}}$: Perplexity exhibits a better generalization capability than feature- and pre-training-based methods. The AUC of supervised methods drops significantly when training and testing sets are generated by different LLMs.
6. Implications
This section discusses the characteristics of LLMgCode and discusses the implications of using Perplexity to detect LLMgCode.
6.1. LLMgCode in Retrospect
The retrospective analysis of LLMgCode primarily focuses on its code length and the resulting impact on the evaluation results.
LLMgCode solutions are inherently more concise compared to HaCode solutions for identical tasks. The LLMgCode solutions and HaCode solutions in our dataset are used to solve the same programming puzzles. However, our dataset has 4,135 HaCode solutions and 8,725 LLMgCode solutions with fewer than 30 LOC. This indicates that LLMgCode solutions are inherently more concise than HaCode solutions. The conciseness of LLMgCode solutions can be explained by three main factors: (1) LLMs are familiar with built-in functions and utility libraries, and prefer using them to simplify implementations, (2) LLMs are trained on large-scale code corpora, and tend to imitate standard and concise coding styles, and (3) LLMs prefer minimal implementations rather than verbose or explanatory styles due to the limitation of tokens.
The conciseness of LLMgCode has a negative impact on the detection accuracy of Perplexity. Perplexity is unsuitable for detecting short code snippets because of its detection mechanism. Since the first few tokens of a code snippet typically have relatively high perplexity due to limited contexts, short solutions are more influenced by this effect than long solutions and therefore tend to have higher perplexity. As a result, Perplexity may misclassify short LLMgCode solutions as HaCode solutions due to their relatively high perplexity. The results of $\text{RQ}_{1.3}$ show that Perplexity does not perform well on small-scale code with an AUC of only 69.75%. As Perplexity is unsuitable for detecting short code snippets, the conciseness of LLMgCode solutions increases the difficulty of detection and reduces the detection accuracy of Perplexity.
To mitigate the impact of code length difference on evaluation results, we evaluate the detection methods from the aspect of code scale in RQ3. Specifically, we evaluate the detection methods on three different code scales, where LLMgCode and HaCode solutions within each scale have comparable lengths. In other evaluation aspects, we do not apply any intervention to limit the code length because the differences in code length arise from the inherent characteristics of code written by LLMs and humans, rather than from our designs.
6.2. For Researchers
Our implications for researchers include both the pros and cons of Perplexity.
@@ 6.2. For Researchers
Perplexity is unsuitable for high-level programming languages. Perplexity performs well in low-level languages such as C/C++ (91%) but poorly in high-level languages such as Python (64%) and Ruby (73%), which implies that Perplexity has limited capability for high-level languages. High-level languages encourage a concise and clear code style. High-level languages such as Python provide many built-in operations for basic elements, e.g., list and dictionary; low-level languages have to manually manage array, pointer, etc. The programming patterns of high-level languages can be more easily learned by a code generator than those of low-level languages. Moreover, high-level languages have more available third-party libraries than low-level languages. Developers are used to leveraging third-party libraries in their daily development to simplify implementation. Since both humans and LLMs use third-party libraries in generating code, detecting LLMgCode from HaCode using Perplexity in high-level languages is challenging.
@@
Perplexity is unsuitable for just-in-time detection. NPR and DetectCodeGPT are the best-performing methods in Perplexity. The detection speeds of NPR and DetectCodeGPT are all around 10 seconds per sample, while the detection speed of the pre-training-based method is less than 0.02 seconds per sample. The difference implies that Perplexity is not suitable for just-in-time detection. This can be explained that Perplexity uses an LLM to calculate the perplexity of each token. The efficiency of LLMs in calculating perplexity for each token is lower than that of pre-trained models in calculating a probability for an entire code snippet. Moreover, NPR and DetectCodeGPT are perturbation-based methods. Such methods calculate the perplexity for one original sample and a collection of perturbed samples. As the number of perturbed samples increases, perturbation-based methods become time-consuming.
@@
Although our investigation indicates that Perplexity has several shortcomings, it exhibits advantages in the following aspects.
@@
Perplexity exhibits the overall good generalization capability. Perplexity with perturbation achieves stable performance in detecting LLMgCode across four LLMs, with AUC variations of around 2%. This implies that Perplexity has a good generalization capability when detecting LLMgCode generated by different LLMs. However, Perplexity without perturbation does not achieve such a performance, e.g., the AUC of LRR varies by more than 10% between LLMs. This highlights the importance of incorporating perturbation when using Perplexity to ensure the generalization capability. The detection mechanism of Perplexity is similar to LLMs for code generation. Therefore, Perplexity remains useful provided that LLMs are based on the auto-regressive generation mechanism (i.e., generating a token based on tokens generated before). The performances of supervised methods are not well when the code snippets for training and testing are generated by different LLMs. The results indicate that supervised methods primarily learn the coding style of LLMs that generate training samples rather than the intrinsic differences between LLMgCode and HaCode. Considering new powerful LLMs are emerging, e.g., DeepSeek, Perplexity is the most suitable method to address this issue.
Perplexity performs well in detecting LLMgCode in C/C++. Perplexity achieves an AUC of 90.40% in C and 92.10% in C++, respectively. This performance is impressive given that Perplexity is a heuristic method rather than a supervised method. These results demonstrate that Perplexity is suitable for detecting LLMgCode in C/C++. C/C++ are low-level programming languages with fewer available third-party libraries than high-level languages such as Python. Therefore, developers implement basic functionality manually. Additionally, C/C++ emphasizes more manual control than high-level languages. Therefore, developers handle details like variable types and dynamic memory. These characteristics make C/C++ code diverse due to personal coding styles and hard for LLMs to imitate. Perplexity performs not as well as the pre-training method in C/C++ LLMgCode of GPT-3.5. However, the pre-training method has poor generalization capability since it is supervised, with an AUC dropping from 92. 21% on GPT-3.5 to 75.21% on Gemini-1.0. In contrast, Perplexity demonstrates the best generalization with the AUC around 80% on all four LLMs, indicating that Perplexity is LLM-independent. Therefore, considering both performance and generalization, we recommend using Perplexity to detect LLMgCode in C/C++.
Perplexity exhibits the good interpretability. Perplexity simulates the code generation mechanisms of LLMs to calculate the perplexity of code snippets. Therefore, the perplexity value is interpretable. A code snippet with a high perplexity indicates that it is disjointed, which might not be LLMgCode. The threshold of perplexity for detection can be adjusted according to the need. Moreover, the calculation of perplexity is performed token by token, allowing it to be precise for the line-level or even token-level detection. This enables Perplexity to perform fine-grained detection. Fine-grained detection does not mean detecting short code. It refers to detecting code snippets in which only a few lines are generated by LLMs. The detection mechanism makes Perplexity applicable to fine-grained detection. For instance, if most of a code snippet is written by a developer and only a few lines are generated by the LLM, Perplexity can calculate the perplexity line by line and detect the lines with low perplexity as LLMgCode. The mechanisms of Perplexity help reviewers understand detection results and provide a foundation for detecting intricate cases where LLMgCode and HaCode are mixed. The mechanisms of supervised methods are complex and hard to interpret. They only output a prediction score or binary result, making developers difficult to understand why the code is detected as LLMgCode and which part is LLMgCode.
In short, there is a trade-off between detection accuracy, speed, and generalization capability. Perplexity is not the most effective in terms of accuracy or efficiency, but it offers the best generalization capability. Unlike supervised methods that learn the characteristics of a specific LLMgCode (e.g., GPT 3.5-generated code), Perplexity seeks to understand the intrinsic differences between LLMgCode and HaCode, making it robust when applied to code generated by unknown LLMs. Although Perplexity is LLM-independent and achieves similar performance across different sources, it has suboptimal accuracy and slower detection speed than feature-based and pre-training-based methods. The best-performing methods of Perplexity are based on perturbation, which introduces additional computational overhead and slows the detection. Therefore, Perplexity is not suitable for just-in-time detection. Separately, Perplexity can be naturally extended to fine-grained detection (e.g., line level) when LLMgCode and HaCode are mixed, whereas other methods provide only a binary classification result (or a prediction score) for the entire code snippet.
Based on the experimental results and the analysis of Perplexity’s detection mechanism, we provide the following directions to improve the detection accuracy and speed of Perplexity.
Effective perplexity calculation. The current detection LLMs used for perplexity calculation have only a few billion parameters, while LLMs used for code generation in practice typically have hundreds of billions of parameters. The large difference in parameters results in significant differences in code understanding and generation, as evidenced by the considerable differences in code solutions generated for the same task. The discrepancy in code understanding and generation between the detection LLM and the generation LLM is a significant reason for the limited effectiveness of Perplexity. To address this problem, first, we recommend using larger LLMs to calculate perplexity, which reduces the differences in perplexity understanding between detection LLMs and generation LLMs, but also increases the computational cost. Second, we recommend distilling the domain knowledge of LLMs with large parameters into LLMs with small parameters. For instance, we can first collect the code solution generated by LLMs with large parameters, e.g., GPT-4o, then utilize fine-tuning or reinforcement learning techniques to train an LLM with small parameters, e.g., CodeLlama-7B, in the domain of code generation, and use the distilled LLMs to calculate the perplexity to detect LLMgCode. By aligning code understanding between detection and generation LLMs, the effectiveness of Perplexity can be improved.
Optimized perturbation method. The perturbation method based on code completion is highly inefficient due to dozens of code masking and completion. Therefore, reducing the number of perturbations is necessary. For instance, instead of randomly masking and completing code lines dozens of times, we can first calculate the perplexity of each line in the original code snippet, and then apply only a few perturbations to the lines with high perplexity. This measure not only improves efficiency several-fold but also amplifies the perplexity discrepancies before and after perturbation, which improves the detection effectiveness. Compared with code completion, adding spaces and empty lines to code snippets is more efficient. However, these perturbation methods are not effective for all types of code, e.g., Python. For indentation-sensitive languages like Python, inserting spaces can cause syntax errors, making the code difficult to understand, i.e., leading to high perplexity for both LLMgCode and HaCode. Therefore, we can further optimize perturbation techniques by considering the characteristics of programming languages. For instance, we can first identify the programming languages of the detected code. If the code is identified as Python, we can apply more targeted perturbation methods, such as replacing variable names. The performance of Perplexity can be improved by implementing optimized perturbation methods.
Improved perplexity utilization. The current utilization of perplexity is insufficient. Representing the perplexity of a code snippet by the average perplexity of each token leads to significant information loss. Two snippets with the same average perplexity could differ in that one has an overall higher perplexity, while the other has a lower overall perplexity but very high perplexity in certain tokens. The perplexity of a code snippet forms a continuous distribution based on the perplexity of each token in sequence. Instead of calculating the average perplexity of each token, we need to better leverage the entire perplexity sequence. For example, we can count the occurrences where the perplexity exceeds a certain threshold, referred to as surpassing occurrences. HaCode tends to be more random and therefore is likely to have more very high perplexity on certain tokens compared to LLMgCode. For another example, we can focus on analyzing changes in the perplexity sequence at the perturbed lines before and after the perturbation. This method amplifies the effect of perturbation, preventing it from being neutralized by the average. The larger the change, the more likely it is LLMgCode, as LLMgCode is more sensitive to perturbations (Shi et al., 2025). Additionally, we can use the perplexity sequence as the initial input for a supervised classifier. Combining the perplexity sequence with supervised learning can improve detection performance compared with Perplexity (heuristic methods).
We propose the following recommendations for the use of Perplexity.
  • Detection method selection:
    • We recommend using DetectCodeGPT or NPR for detection under normal circumstances, as they show the best detection accuracy, acceptable detection speed, and good generalization capability. DetectCodeGPT and NPR perform best among Perplexity, achieving an AUC of around 80% for different LLMs. Moreover, DetectCodeGPT and NPR demonstrate an acceptable detection speed, detecting a sample in just a few seconds.
    • We recommend using Log-Rank for detecting large volumes of samples or for just-in-time detection, as it offers acceptable effectiveness and a fast detection speed. Although Perplexity with perturbation offers better detection performance, it takes several tens of seconds to detect a single sample. Log-Rank performs best among Perplexity without perturbation, with an AUC close to 70% in most cases, while taking only a fraction of a second to detect a sample (50 times faster than DetectCodeGPT/NPR).
  • Application scenario of Perplexity:
    • We recommend applying Perplexity to detect LLMgCode in C/C++ or with a large scale. For the programming language, Perplexity exhibits good performance in C/C++, with an AUC above 90%, while performing suboptimal in other languages like Python, where it only reaches an AUC of 63%. For the code scale, Perplexity exhibits good performance in code solutions with a large scale, achieving an AUC of 87%, while it has limited effectiveness in code solutions with a small or middle scale.
    • We recommend applying Perplexity when the generation models of LLMgCode are unknown, as Perplexity has the best generalization capability. Perplexity achieves an AUC of around 80% regardless of LLMs. However, supervised methods learn the coding styles of LLMs that generate training samples, leading to significant performance fluctuations across different LLMs. Given the continuous emergence of new LLMs and the availability of numerous LLMs for code generation, it is difficult to know whether the training set includes samples generated by the model that generated the code to be detected.
    • We recommend applying Perplexity for fine-grained detection since it can be applied to line-level detection. When a code snippet contains both HaCode and LLMgCode, feature- and pre-training-based methods can only assign a single predictive score to the entire snippet and can not make accurate distinctions. In contrast, Perplexity calculates the perplexity token by token, allowing it to assess the perplexity of each line individually.
6.3. For Tool Development and Usages